Compute the change in internal energy when 50g of water is turned into 8.4 x 104 cm³ of steam at 100°c under normal Atm. pressure. The latent heat of vaporization of water is 2.26 x 106 J/kg.
1
Expert's answer
2020-12-10T11:10:15-0500
"\Delta U=Q+W=\Delta H m-p\Delta V\\
\Delta U=-(1)(84-0.05)(101.3)+2.26\cdot10^{6}(0.05)\\\Delta U=104500\ J"